# Leveraging Multimodal LLM for Inspirational User Interface Search

### v2

@@ 7.2. Enhancing the Design Inspiration Process
Our user study findings indicate that semantic-based search significantly enhances designers’ ability to find relevant inspirational screens. Designers reported that the system’s semantic filters allowed them to perform more granular searches, focusing on specific screen roles, user flows, or target audiences. For example, a designer looking for “onboarding screens for a fitness app targeting beginners” could retrieve highly relevant examples by searching semantics (screen role: onboarding, app category: health and fitness, target user: fitness beginner), which would be difficult to find using traditional keyword searches.
@@
Providing MLLM’s semantic explanations alongside search results increased designers’ trust in the system and made the search process more transparent. Designers appreciated seeing the extracted semantics for each UI screen, which allowed them to understand why specific results were retrieved. This transparency enabled them to decide which designs to consider further. One participant mentioned, “Seeing the semantic explanation helps me quickly assess if a design is relevant to my project. It feels like the system understands what I’m looking for.” This feedback highlights the importance of explainability in design tools, as it enhances usability and user satisfaction.
@@ 7.3. Implications for UI Design Search Tools
Our findings suggest integrating functional and aesthetic support in design tools aligns well with designers’ needs. Designers often seek inspiration that fulfills specific functional purposes within an app while also being visually appealing. By providing semantic filters for both aspects, the system helps designers find relevant and aesthetic designs. For instance, a designer working on a financial app can search for “transaction history screens with a minimalist design,” ensuring that the retrieved examples meet both the functional requirements and the desired aesthetic style. This multi-faceted support facilitates the design process by consolidating search efforts into a single tool.
@@
Providing semantic analysis results for each screen aids designers in iterative exploration. Designers can refine their searches based on the extracted semantics of previously viewed screens, leading to a more efficient and targeted exploration process. This iterative approach enables designers to explore broader possibilities while staying aligned with their project goals. Moreover, the semantic information facilitates better communication between design teams and other stakeholders. By having a set of semantic descriptors, team members can discuss design elements more effectively. For example, referring to a “joyful, booking and reserving flow for active people” provides a clear and shared understanding of the design context, reducing misunderstandings and improving collaboration.
@@ 7.4. Public Release of the S&UI Dataset
As part of our commitment to advancing research in UI design, we are publicly releasing the S&UI Dataset used in this study’s experiments on GitHub [^3] https://github.com/spark-damian/S-UI. This dataset includes semantic annotations extracted from the UI screenshots used in our evaluations. By making this dataset available, we aim to provide the research community with a practical resource that can be used to replicate our findings and explore new approaches to UI semantic analysis. This release of S&UI dataset represents a significant step towards fostering collaboration and innovation in UI design, enabling others to build upon our work and drive new developments in semantic-based UI retrieval, flow analysis, and design comparison.
@@ 7.5. Limitation and Future Directions
While our studies demonstrated promising results, we acknowledge limitations and areas for future research. First, our current query interface requires manual input of semantic criteria, which may require additional thought, especially during the early stages of the design process when designers are often exploring broad ideas rather than specific attributes. Future work should explore allowing designers to express their needs through free-form queries (e.g., “Show me a minimalist login screen with dark colors”), automatically mapping these to each semantic attribute. Second, our retrieval method’s reliance on weighted embedding similarity becomes less effective as query complexity increases. Alternative approaches like semantic combination filtering or category-based matching could provide more relevant and diverse results. Also, our current implementation focuses on individual screens, while designers often need to analyze larger patterns across multiple screens and applications. Expanding to multi-screen and cross-app analysis could better support user flow evaluation, including screen transitions, interaction patterns, and design consistency. Likewise, while our work primarily focuses on mobile UI design, the underlying approaches and techniques could be adapted and generalized to other design contexts, such as web pages or desktop applications. Adapting to these contexts would likely involve diversifying the semantic extraction pipeline to effectively manage the increased complexity of layouts and the broader range of elements. Another exciting direction is exploring how MLLMs could enable new semantic-aware design synthesis and exploration capabilities in generative design tools. By pursuing these directions, we can work towards more capable, flexible, and intuitive tools that support designers’ complex needs and inspire new forms of semantic-aware design exploration.

### v3

@@ 7.2. Enhancing the Design Inspiration Process
Our user study findings indicate that semantic-based search significantly enhances designers’ ability to find relevant inspirational screens. Designers reported that the system’s semantic filters allowed them to perform more granular searches, focusing on specific screen roles, user flows, or target audiences. For example, a designer looking for ”onboarding screens for a fitness app targeting beginners” could retrieve highly relevant examples by searching semantics (screen role: onboarding, app category: health and fitness, target user: fitness beginner), which would be difficult to find using traditional keyword searches.
@@
Providing MLLM’s semantic explanations alongside search results increased designers’ trust in the system and made the search process more transparent. Designers appreciated seeing the extracted semantics for each UI screen, which allowed them to understand why specific results were retrieved. This transparency enabled them to decide which designs to consider further. One participant mentioned, ”Seeing the semantic explanation helps me quickly assess if a design is relevant to my project. It feels like the system understands what I’m looking for.” This feedback highlights the importance of explainability in design tools, as it enhances usability and user satisfaction.
@@ 7.3. Implications for UI Design Search Tools
Our findings suggest integrating functional and aesthetic support in design tools aligns well with designers’ needs. Designers often seek inspiration that fulfills specific functional purposes within an app while also being visually appealing. By providing semantic filters for both aspects, the system helps designers find relevant and aesthetic designs. For instance, a designer working on a financial app can search for ”transaction history screens with a minimalist design,” ensuring that the retrieved examples meet both the functional requirements and the desired aesthetic style. This multi-faceted support facilitates the design process by consolidating search efforts into a single tool.
@@
Providing semantic analysis results for each screen aids designers in iterative exploration. Designers can refine their searches based on the extracted semantics of previously viewed screens, leading to a more efficient and targeted exploration process. This iterative approach enables designers to explore broader possibilities while staying aligned with their project goals. Moreover, the semantic information facilitates better communication between design teams and other stakeholders. By having a set of semantic descriptors, team members can discuss design elements more effectively. For example, referring to a ”joyful, booking and reserving flow for active people” provides a clear and shared understanding of the design context, reducing misunderstandings and improving collaboration.
@@ 7.4. Public Release of the S&UI Dataset
As part of our commitment to advancing research in UI design, we are publicly releasing the S&UI Dataset used in this study’s experiments on GitHub [^3] https://github.com/spark-damian/S-UI. This dataset includes semantic annotations extracted from the UI screenshots used in our evaluations. By making this dataset available, we aim to provide the research community with a practical resource that can be used to explore new approaches to UI semantic analysis. This release of S&UI dataset represents a significant step towards fostering collaboration and innovation in UI design, enabling others to build upon our work and drive new developments in semantic-based UI retrieval, flow analysis, and design comparison.
@@ 7.5. Limitation and Future Directions
While our studies demonstrated promising results, we acknowledge limitations and areas for future research. First, our current query design requires manual input of each semantics, which may require additional thought, especially during the early stages of design processes where designers often explore broad ideas rather than specific attributes. Future work should explore allowing designers to express their needs through free-form queries (e.g., ”Show me a minimalist login screen with dark colors”), automatically mapping these to each semantic attribute. Second, our retrieval method’s reliance on weighted embedding similarity becomes less effective as query complexity increases. Alternative approaches, such as semantic combination filtering or category-based matching, could provide more relevant and diverse results. Also, our current implementation focuses on individual screens, while designers often need to analyze larger patterns across multiple screens and applications. Expanding to multi-screen and cross-app analysis could better support user flow evaluation, including screen transitions, interaction patterns, and design consistency. Likewise, while our work focuses on mobile UI design, its approaches and techniques could extend to web and desktop applications by diversifying the semantic extraction pipeline to handle more complex layouts and elements. Another exciting direction is exploring how MLLMs could enable new semantic-aware design synthesis and exploration capabilities in generative design tools. Additionally, a deeper analysis of how expertise shapes AI-driven design search workflows and use cases could provide valuable insights. By pursuing these directions, we can work towards more capable, flexible, and intuitive tools that support designers’ complex needs and inspire new forms of semantic-aware design exploration.